A boy standing on a long railroad car throws a ball straight upwards. The car is moving on the horizontal road with an acceleration of 1 m/s2 and the projection velocity in the vertical direction is 9.8 m/s. How far behind the boy will the ball fall on the car?

Grade:upto college level

1 Answers

Navjyot Kalra
9 years ago
Sol. Let the velocity of car be u when the ball is thrown. Initial velocity of car is = Horizontal velocity of ball. Distance travelled by ball B Sb = ut (in horizontal direction) And by car Sc = ut + 1/2 at2 where t → time of flight of ball in air. ∴ Car has travelled extra distance Sc – Sb = 1/2 at2. Ball can be considered as a projectile having θ = 90°. ∴ t = (2u sinθ)/g = (2 x 9.8)/9.8 = 2 sec. ∴ Sc – Sb = 1/2 at2 = 2 m ∴ The ball will drop 2m behind the boy.